OpenStudy (megannicole51):

By setting one variable constant, find a plane that intersects the graph of z=(x^2+1)sin(y)+xy^2 in a: (a) parabola (b) straight line (c) sine curve

ganeshie8 (ganeshie8):

@ikram002p

OpenStudy (anonymous):

try setting x= a constant and see what you get, then see what you would have to set x to to get a sine curve. then try setting y= a constant and see what you get

OpenStudy (megannicole51):

x=2 5sin(2)+2y^2 4.5+2y^2

OpenStudy (megannicole51):

idk what u mean by a sine curve

OpenStudy (megannicole51):

im sorry i typed that in wrong

OpenStudy (megannicole51):

x=2 5sin(y)+2y^2

OpenStudy (helder_edwin):

if u put x=0 then \[\large z=(x^2+1)\sin y+xy ^2 \] becomes \[\large z=(0+1)\sin y+0\cdot y^2=\sin y \] which is a sine curve

OpenStudy (helder_edwin):

see what u get with \[\large y=\pi/2 \] and \[\large y=\pi \]

OpenStudy (megannicole51):

when y=pi/2 x^2+1+x*(pi^2/4)

OpenStudy (megannicole51):

when y=pi xpi^2

OpenStudy (megannicole51):

thank you so much for your help!!

OpenStudy (helder_edwin):

right. if \(y=\pi/2\) then \[\large z=(x^2+1)\sin y+xy^2 \] becomes \[\large z=(x^2+1)\sin(\pi/2)+x(\pi/2)^2 \] \[\large z=(x^2+1)\cdot1+x(\pi^2/4) \] \[\large z=x^2+1+(\pi^2/4)x \] \[\large z-1=x^2+(\pi^2/4)x+(\pi^2/8)^2 \] \[\large z-1=(x+\pi^2/8)^2 \] is the parabola on the xz-plane with vertix \((-\pi^2/8,0,1)\).

OpenStudy (helder_edwin):

if \(y=\pi\) then \[\large z=(x^2+1)\sin\pi+x\pi^2 \] \[\large z=(x^2+1)\cdot0+x\pi^2 \] \[\large z=\pi^2x \] which is not a line, it is actually a plane (remember u r in R^3). [I made a mistake]. I don't think u can get a line from that equation (at least i don't see how)